### v2

@@
Imagine you and a friend purchase identical items at a store, yet only your friend received a discount. Would your friend’s discount make you feel unfairly treated by the store? And would you be less willing to purchase from that store again in the future? Based on a large-scale online survey that we ran on Prolific, it turns out that the answers to the above questions are positive. Motivated by these findings, in this work we propose a notion of individual fairness in online revenue management and an algorithmic module (called “Grace Period”) that can be embedded in traditional revenue management algorithms and guarantee individual fairness. Specifically, we show how to embed the Grace Period in five common revenue management algorithms including Deterministic Linear Programming with Probabilistic Assignment (DLP-PA), Resolving Deterministic Linear Programming with Probabilistic Assignment (RDLP-PA), Static Bid Price Control (S-BPC), Booking Limit (BL), and Nesting (NESTING), thus covering both stochastic and adversarial customer arrival settings. Embedding the Grace Period does not incur additional regret for any of these algorithms. This finding indicates that there is no tradeoff between a seller maximizing their revenue and guaranteeing that each customer feels fairly treated.
@@ 1 Introduction
This work emphasizes the equitable treatment of customers through the requirement of individual fairness. The concept of fairness is crucial not only from an ethical standpoint but also for its impact on customer behavior and potential to increase revenue. For example, think about the following (hypothetical, for now) scenario. Two stores, A and B, offer an identical product at a same price. However, Store A selectively offers discounts leading to perceived unfairness among customers. One may think that such perceived unfair treatment may deter customers from buying at Store A, possibly benefiting Store B. Okun (2011) explains that fair treatment to each individual is instrumental for a profit-maximizing firm to the maximization of long-run profits. To confirm if customer reactions align with our aforementioned intuition, we conducted a large-scale online survey on Prolific, detailed in Section 7. Our survey’s findings validate our priors; customers do indeed perceive various forms of unfair treatment from sellers and this in turn affects the probability of them buying again from certain stores.
We next define individual fairness in the context of revenue management.
Definition 1.1 (Individual Fairness in Revenue Management)
Let $n$ be the number of customer types, with each type $i\in[n]$ having $n_{i}$ customers arriving. An algorithm $\mathcal{A}$ is $(\alpha,\delta)$-fair for any choice of $\alpha,\delta\in(0,1)$ if for each customer of type $i$, with a probability at least $1-\delta$, for any $u,v\in[n_{i}]$,
$\mathbb{P}\left[i(v)\succ_{\mathcal{A}}i(u)\right]\leq\alpha|u-v|,$
where $i(u),i(v)$ denote the $u$-th and $v$-th customer’s of type $i$, and $i(v)\succ_{\mathcal{A}}i(u)$ denotes that algorithm $\mathcal{A}$ accepts customer $i(v)$ but not $i(u)$.
Intuitively speaking, an algorithm satisfying individual fairness ensures that the probability of two customers of the same type that arrive closely in time but receive disparate treatments is very small. Our definition of individual fairness (both the time- and the type-dependent components) is driven by our survey findings (cf. Section 7), which show that people often perceive unfairness when they do not obtain a resource that other people that are similar to them receive. As a result, we aim to upper bound the probability of customers of the same type ($i(v)$ and $i(u)$) receiving disparate treatment. As for the time component, Kahneman et al. (1986) executed a telephone survey inquiring whether customers perceived differential treatment by a hardware store during a night and the subsequent early morning, amidst a severe snowstorm, as fair. The findings indicated that 82% of respondents deemed such disparate treatment as unfair. Furthermore, our survey findings suggest that individuals perceive different treatments occurring in the near term as more unfair compared to those in the distant past. This insight suggests that the probability of two similar customers being treated differently, $\mathbb{P}\left[i(v)\succ_{\mathcal{A}}i(u)\right]$, should be moderated according to the time gap between their arrivals, specifically using the L1 norm, $\alpha|u-v|$, as a measure. Furthermore, our findings remain valid when considering measures that increase linearly or at a greater rate with respect to $|u-v|$, i.e., $\alpha|u-v|^{p}$ for any $p\geq 1$. In the existing literature, this measure is referred to as the time-variant individual fairness metric, as suggested by Gupta and Kamble (2021). A comprehensive discussion of various fairness metrics presented in prior research is available in Section 1.2.
@@
In this paper, our primary objective is to incorporate the principle of individual fairness into the most commonly used algorithms solving the quantity-based revenue management problem, including Deterministic Linear Programming with Probabilistic Assignment (DLP-PA), Resolving Deterministic Linear Programming with Probabilistic Assignment (RDLP-PA), Static Bid Price Control (S-BPC), Booking Limit (BL), and Nesting (NESTING). We modify each algorithm one by one to meet individual fairness criteria as defined in Definition 1.1. In addition to implementing individual fairness, we also strive to maintain revenue optimization, ensuring that incorporating fairness does not lead to any loss compared to the original algorithm’s performance [^1] In Appendix 13, we show how our approach can also produce an individually-fair variant of Bid Price Control with Online Gradient Descent (BPC-OGD), albeit with a small asymptotic loss in revenue.. By adapting these well-known algorithms, we aim to ensure that any revenue management problems addressed by them also upholds the principles of individual fairness.
@@ 1.1 Main Contributions
“Grace Period” Design. In Section 3, we introduce the Grace Period which we find to be the workhorse concept for achieving individual fairness in revenue management. At a simplified level, the Grace Period works as follows: As the seller’s available resources approach depletion, an arriving customer is either accepted or rejected based on the treatment of the preceding customer. If the previous customer was accepted, the succeeding customer is likewise accepted with a probability of $(1-\alpha)$. However, if the previous customer was rejected, the incoming customer is also declined. This design (which gets utilized only when the resources are almost depleted) forms a grace period during which the probability of unequal treatment is confined. We remark that the Grace Period for customers of type $i$ can (and will in general) be different from the Grace Period for customers of type $j$ (for $i\neq j$). The tricky part is to guarantee that despite these distinct grace periods, the seller does not sell resources beyond their allowable capacity. We showcase how the Grace Period can be used via an example analysis of the First-Come-First-Serve (FCFS) algorithm.
@@
Grace-Period-Enhanced Algorithms for Revenue Management. In Sections 4 and 5, we integrate the Grace Period design into five well-known algorithms frequently used in revenue management (RM): DLP-PA, RDLP-PA, S-BPC, BL and NESTING. The first three DLP-PA, RDLP-PA, S-BPC are used for stochastic customer arrivals, while BL and NESTING are used for adversarial arrivals. For completeness, we outline each of these algorithms together with the explanation of why they are not individually fair in Section 2. Based on the different structure of these algorithms, we employ different methods to incorporate the grace period design into them. For example, in DLP-PA, each type $i$ customer is accepted with certain probability $p_{i}$, leading to unfairness because consecutive type $i$ customers are treated differently with a probability $p_{i}(1-p_{i})$, violating our fairness definition (1.1) for $\alpha<p_{i}(1-p_{i})$. To address this, we avoid random decisions for individual arrivals. Instead, we divide the timeline into segments, accepting a random number of customers in a first-come-first-serve manner per segment. Grace periods inside each segment ensure fairness when decisions change from acceptance to rejection within each segment; grace periods between each segment guarantee fairness when decisions change from rejection to acceptance between the segments. For clarity, we refer to the new algorithms as GP-Enhanced.
@@
No Trade-off between Revenue Maximization and Individual Fairness in an Asymptotic Regime. Contrary to the common belief that there exists a trade-off between maximizing revenue and achieving individual fairness, we demonstrate that by augmenting standard RM algorithms with the grace period design, it is possible to attain optimal revenue (asymptotically) while simultaneously ensuring individual fairness for most of the algorithms.
@@
For settings where the customer arrival process is stochastic, the performance of RM algorithms is measured in terms of regret with regards to the benchmark of the optimal revenue achievable (which assumes known realized demand). Table 1 contains a summary of the regret for GP-Enhanced algorithms for stochastic arrivals versus the traditional non-individually fair counterparts. For settings where the customer arrival process is adversarial, the performance of RM algorithms is measured in terms of the competitive ratio; that is, the worst-case ratio (across all possible arrival instances) between the revenue generated by the algorithm and the hindsight optimal (i.e., the maximum revenue assuming full knowledge of the realized demand). Table 2 contains a summary of the competitive ratio for GP-Enhanced algorithms for adversarial arrivals versus the traditional non-individually fair counterparts.
@@
Table 1: Regret of Original Algorithms and Grace Period Enhanced Algorithms under Stochastic Arrival
Original Algorithm Regret of Original Algorithm Regret of Grace Period Enhanced Algorithm
@@
RDLP-PA $O(T^{1/3})$ $O(T^{1/3})$ (Theorem 4.1)
@@
Table 2: Competitive Ratio of Original Algorithms and Grace Period Enhanced Algorithms under Adversarial Arrival. $C\in[0,1]$ is a setting-specific value that is specified in the analysis.
Original Algorithm Competitive Ratio of Original Algorithm Competitive Ratio of Grace Period Enhanced Algorithm
BL $C$ $C-O(\log m/m)$ (Theorem 5.1)
NESTING $C$ $C-O(\log m/m)$ (Theorem 5.2)
@@
No Additional Computational Cost for Implementing Individual Fairness. One might conjecture that simultaneously achieving optimal revenue and being individually fair might necessitate additional computational cost for the algorithm. However, this is not the case. In all GP-Enhanced algorithms, we have preserved the algorithms’ original structure to the maximum extent possible, maintaining the order of computational complexity equivalent to that of the original algorithm.
@@
Survey Highlighting the Significance of Individual Fairness in RM. We conducted an online survey between October and November 2023 on Prolific about people’s perception of unfair treatment in RM. Our findings suggest: (i) When customers witness behavior in a store that is considered “unfair” (as described in Definition 1.1) before they shop, their inclination to buy decreases. (ii) A customer who has bought from a store in the past and then experiences what they perceive as unfair treatment is likely to decrease their purchasing from that store in the future. (iii) Individuals perceive different treatment happening in the near term as more unfair compared to those in the distant past. With these observations, we can evaluate the importance of individual fairness to business and validate the mathematical definition of individual fairness metrics in Definition 1.1.
1.2 Other Related Work
Time Varient Individual Fairness. In the fields of Computer Science and Operations Research, individual fairness, often referenced as envy-freeness, is a critical metric in scenarios where equitable treatment of entities—such as customers, users, or resources—is desired. The time-variant aspect of individual fairness adds another layer of complexity, as it demands that customers arriving at similar times should receive similar treatment. Different papers adapt the concept of fairness to suit their specific context, which can lead to varying definitions and formulations. Here are some general axioms and mathematical formulations that capture the essence of time-variant individual fairness:
Axiom 1: Lipschitz Continuity in Time (Dwork et al. 2012, Gupta and Kamble 2021, Mukherjee et al. 2020)
This formulation states that the change in treatment should be limited by the elapsed time between arrivals, ensuring a smooth transition:
$D(f(t),f(s))\leq Ld(t,s)$
Here, $t$ and $ts$ represent the arrival time, $f(x)$ denotes the treatment function, $D$ is a distance metric measuring similarity between treatments, $d$ is a distance metric measuring the time difference, and $L$ is a Lipschitz constant that bounds the rate of change in treatment over time. This ensures that the treatment function $f$ does not vary too sharply with small changes in time.
Axiom 2: Probabilistic Fairness (Our paper), (Donahue and Kleinberg 2020, Herlihy et al. 2023)
This approach introduces randomness in treatment while maintaining fairness statistically:
$P(D\left(f(t),f(s)\right)|\leq\delta|t,s)\geq 1-\epsilon\text{ for }d(t,s)<\gamma$
where $\delta,\gamma,\epsilon$ are thresholds dictating the acceptable level of treatment difference, time difference, and probability of unequal treatment, respectively.
Axiom 3: Time-Consistent Treatment (Cohen et al. 2021, Gupta et al. 2023, Heidari and Krause 2018)
This axiom ensures that entities arriving within a short time window receive comparable treatment. Mathematically, it can be expressed as:
$\forall t,s\text{ with }d(t,s)<\epsilon,\quad D(f(t),f(s))\leq\delta$
Here, $\epsilon,\delta$ are small constants defining the permissible time window and treatment disparity, respectively.
Our fairness metric aligns with Axiom 2. Given that our time horizon is discrete, we adopt the distance measure $d(t,s)=|t-s|$. Furthermore, we have verified that substituting this with $d(t,s)=|t-s|^{p}$ for any $p\geq 1$ does not affect the validity of our results. Moreover, unlike the literature where decision variables are continuous—such as pricing or the allocation of divisible goods—our decision variable is discrete, limited to either accept or reject. Consequently, we employ $D$ as a probability measure to assess the likelihood of different treatments, which is the most appropriate measure in our context.
Mathematically speaking, Axiom 1 is stronger than Axiom 2. The intuition behind is that in Axiom 1, the treatment function must not change more rapidly than a constant rate over time. This means that the rate at which treatment changes is tightly controlled, providing a uniform smoothness in how treatments are administered over time.
Axiom 2 is stronger than Axiom 3. The intuition is that Axiom 2 generally specifies that for entities arriving within a close timeframe, the probability of receiving similar treatment must be high—usually quantified by a certain threshold. It requires a statistical guarantee that can be measured and verified. Axiom 3, while ensuring that entities arriving at similar times receive similar treatment, doesn’t specify how similar the treatments need to be or provide a statistical probability of similarity. It’s more qualitative and less stringent, making it weaker in comparison.
Individual Fairness from the Regulator’s Perspective. Individual fairness can be categorized from distinct perspectives. The first perspective considers the regulator’s perspective (Arsenis and Kleinberg 2022), suggesting that the acceptance probability for each arriving customer should be equal. The second type looks at fairness from the customer’s perspective, seeking to reduce the likelihood of individuals experiencing envy and subsequently minimize the loss associated with such envy. Our research, along with all works discussed in the previous time-variant individual fairness section, falls under this category.
These two dimensions of individual fairness differ due to their distinctive focuses. Consider an example: two customers arrive, and the first algorithm accepts both customers A and B with a probability of $0.5$. The second algorithm accepts both customers A and B with a probability of $0.1$. It appears that both algorithms uphold individual fairness from the regulator’s perspective. However, the probability that customer 1 is rejected and customer 2 is accepted under the first algorithm is $(1-0.5)\cdot 0.5=0.25$, whereas for the second algorithm, it is $(1-0.1)\cdot 0.1=0.09$. This implies that the second algorithm would potentially lead to lower feelings of envy and unfairness among customers, thus demonstrating the difference in the two metrics of individual fairness.
Modern Revenue Management Models. Due to the limitations of the traditional network revenue management model, researchers have diversified its approach for greater relevance and depth. Despite these advancements, the resolution of these contemporary models often relies on the foundational algorithmic frameworks discussed in Section 2. Our contributions aim to refine these foundational structures, thereby promoting individual fairness across all advanced revenue management models that employ decision-making algorithms outlined in Section 2. For instance, to ensure group fairness, a DLP-PA algorithm is proposed by Ma et al. (2020), while a BPC-OGD algorithm is suggested by Balseiro et al. (2021). In the context of choice-based revenue management, a choice-based DLP-PA algorithm is recommended by Liu and Van Ryzin (2008), Gallego et al. (2015). To tackle revenue management issues involving offline prediction in demand, Balseiro et al. (2022), Golrezaei et al. (2023) advocate for a BL algorithm. Revenue management with correlated arrivals is addressed by Jiang (2023) with an S-BPC method, and Jiang et al. (2020) proposes a BPC-OGD algorithm for non-stationary arrivals. In scenarios involving revenue management with reusable resources, Huo and Cheung (2022) employs a NESTING algorithm. Finally, for revenue management with horizontal uncertainty, a BPC-OGD algorithm is suggested by Balseiro et al. (2023a).
Price-based Revenue Management. Contrasting with the quantity-based model in this paper, price-based revenue management represents a distinct category. Here, pricing decisions inversely affect demand, with higher prices leading to lower demand and vice versa. Maglaras and Meissner (2006) suggests that, under certain conditions, price-based and quantity-based models align in function. Furthermore, price-based strategies also utilize the algorithmic frameworks mentioned in Section 2. Works by Gallego and Van Ryzin (1994, 1997) implement the DLP-PA algorithm for decision-making, whereas Jasin and Kumar (2012), Jasin (2014), Wang and Wang (2022) employ the RDLP-PA algorithm. In Section 8, we explore the concept of individual fairness within the price-based framework and discuss the integration of a grace period design in this context.
Group Fairness. While a considerable volume of literature exists concerning fairness in RM algorithms, much of it primarily aims to achieve group fairness (Cohen et al. 2022, Aminian et al. 2023, Balseiro et al. 2021, Cayci et al. 2020, Donahue and Kleinberg 2020, Ma et al. 2020, Manshadi et al. 2021). Group fairness in decision-making entails ensuring that every group (or type) of customer is treated equitably. However, group fairness does not guarantee individual fairness. This is because, firstly, individuals might not be aware of the groups to which they belong. Secondly, individuals within each group can be treated inequitably based on their specific contexts and features.
2 Model and Preliminaries

### v3

@@
Problem definition: Imagine you and a friend purchase identical items at a store, yet only your friend received a discount. Would your friend’s discount make you feel unfairly treated by the store? And would you be less willing to purchase from that store again in the future? Based on a large-scale online survey that we ran on Prolific, it turns out that the answers to the above questions are positive. Therefore, when allocating resources to different customers, sellers should consider both the total reward and individual fairness.
Methodology/results: Motivated by these findings, in this work we propose a notion of individual fairness in online revenue management and an algorithmic module (called “Grace Period”) that can be embedded in traditional revenue management algorithms and guarantee individual fairness. Specifically, we show how to embed the Grace Period in five common revenue management algorithms including Deterministic Linear Programming with Probabilistic Assignment (DLP-PA), Resolving Deterministic Linear Programming with Probabilistic Assignment (RDLP-PA), Static Bid Price Control (S-BPC), Booking Limit (BL), and Nesting (NESTING), thus covering both stochastic and adversarial customer arrival settings. Embedding the Grace Period does not incur additional regret for any of these algorithms. This finding indicates that, in an asymptotic regime, there is no tradeoff between a seller maximizing their revenue and guaranteeing that each customer feels fairly treated.
Managerial implications: The core intuition behind the Grace Period is that independent randomized decisions for each customer often lead to unfair outcomes. However, we cannot eliminate the randomness, as it plays a crucial role in maximizing profit. The Grace Period addresses this by shifting randomness away from individual decisions and applying it instead to the total number of customers receiving a particular decision. This approach preserves revenue potential while mitigating fairness issues. Furthermore, implementing the Grace Period depends on the specific decision algorithm, with tailored methods provided for each to incorporate this mechanism effectively.
@@ 1 Introduction
This work emphasizes the equitable treatment of customers through the requirement of individual fairness. The concept of fairness is crucial not only from an ethical standpoint but also for its impact on customer behavior and potential to increase revenue. For example, think about the following (hypothetical, for now) scenario. Two stores, A and B, offer an identical product at a same price. However, Store A selectively offers discounts leading to perceived unfairness among customers. One may think that such perceived unfair treatment may deter customers from buying at Store A, possibly benefiting Store B. Xia et al. (2004) support this intuition by showing that when customers perceive a high degree of transaction similarity—such as purchasing the same product under similar conditions or within a short time frame—price discrepancies are more likely to be judged as unfair. They further argue that perceived unfairness can trigger both rational, money-focused responses (e.g., complaints or refund requests) and emotional reactions (e.g., anger or revenge), ultimately eroding customer trust in the seller. Reinforcing this view, Okun (2011) explains that fair treatment of each individual is instrumental for a profit-maximizing firm to the maximization of long-run profits. To support even further that customer reactions align with our aforementioned intuition, we conducted a large-scale online survey on Prolific, detailed in Section 8. Our survey’s findings validate both our priors and the previous research; customers do indeed perceive various forms of unfair treatment from sellers and this in turn affects the probability of them buying again from certain stores.
To incorporate fairness into revenue management, we need a meaningful way to measure the extent of equitable treatment. A fully fair policy—where all customers requesting the same resource are either always accepted or always rejected—would eliminate any unfairness but could significantly reduce revenue by failing to allocate resources efficiently. Instead, a more practical approach is to quantify fairness probabilistically, assessing the likelihood that two customers requesting the same resource receive different decisions. This allows us to evaluate fairness while preserving decision-making flexibility and avoiding rigid rules that could harm revenue optimization.
Building on the idea that fairness should be measured probabilistically, we must carefully consider which probability matters. There are two natural ways to define fairness: one from the decision-maker’s perspective and one from the customer’s perspective. The decision-maker might consider fairness in terms of ensuring that all customers requesting the same resource are accepted with the same probability (see e.g., (Arsenis and Kleinberg 2022)). However, this does not fully capture the individual customer’s perception of fairness, which is shaped by the likelihood that two such customers receive different outcomes—an event that can lead to a rejected customer perceiving unfair treatment. Since fairness concerns arise when one customer is accepted while another is rejected (Xia et al. (2004)), we focus on minimizing the probability of such disparities to ensure that perceived unfairness occurs as infrequently as possible.
To illustrate the difference, consider two algorithms. In the first, two customers requesting the same resource are each accepted with probability 0.3, while in the second, both are accepted with probability 0.5. From the decision-maker’s standpoint, both algorithms appear fair since they treat all such customers identically in terms of individual acceptance probabilities. However, their actual impact on customers differs significantly. In the first case, given that the second customer observes that the first one is accepted, the probability to be rejected is $0.3$, while in the second case, this probability increases to $0.5$. Thus, even when an algorithm assigns equal acceptance probabilities, the probability of observed disparate treatment may vary, meaning fairness at the decision-making level does not necessarily translate to fairness as experienced by customers.
In this paper, we focus on fairness from the customer’s perspective, ensuring that the probability of disparate treatment—where one customer is accepted while another is rejected—is explicitly bounded. The exact definition of the fairness metric is provided in Section 3, where we closely examine its key aspects. We recognize that there is no single, unified way to define fairness; we have included an extensive discussion with our thoughts about how we came up with this definition in Section 10.
@@
In this paper, our primary objective is to incorporate the principle of individual fairness into the most commonly used algorithms solving the quantity-based revenue management problem, including primal methods such as Deterministic Linear Programming with Probabilistic Assignment (DLP-PA), Resolving Deterministic Linear Programming with Probabilistic Assignment (RDLP-PA), dual methods such as Static Bid Price Control (S-BPC), Dynamic Bid Price Control (BPC-OGD), quota-based methods such as Booking Limit (BL), and Nesting (NESTING). We modify each algorithm one by one to meet individual fairness criteria as defined in Definition 3.1. In addition to implementing individual fairness, we also strive for revenue optimization, ensuring that incorporating fairness does not lead to any loss compared to the original algorithm’s performance. By adapting these well-known algorithms, we aim to ensure that any revenue management problems addressed by them also uphold the principles of individual fairness.
@@ 1.1 Main Contributions
“Grace Period” Design. In Section 4, we introduce the Grace Period which we find to be the workhorse concept for achieving individual fairness in revenue management. At a simplified level, the Grace Period works as follows: As the seller’s available resources approach depletion, an arriving customer is either accepted or rejected based on the treatment of the preceding customer. If the previous customer was accepted, the succeeding customer is likewise accepted with a probability of $(1-\alpha)$, where $\alpha$ is a small fairness parameter capturing the allowable probability of unequal treatment; and the rigorous definition can be found in Definition 3.1 in Section 3. However, if the previous customer was rejected, the incoming customer is also declined. This design (which gets utilized only when the resources are almost depleted) forms a grace period during which the probability of unequal treatment is confined. We remark that the Grace Period for customers of type $i$ can (and will in general) be different from the Grace Period for customers of type $j$ (for $i\neq j$). The tricky part is to guarantee that despite these distinct grace periods, the seller does not sell resources beyond their allowable capacity. We showcase how the Grace Period can be used via an example analysis of the First-Come-First-Serve (FCFS) algorithm.
@@
Grace-Period-Enhanced Algorithms for Revenue Management. In Sections 5 and 6, we integrate the Grace Period design into six well-known algorithm structures frequently used in revenue management (RM): DLP-PA, RDLP-PA, S-BPC, BPC-OGD, BL and NESTING. The first four DLP-PA, RDLP-PA, S-BPC, BPC-OGD are used for stochastic customer arrivals, while BL and NESTING are used for adversarial arrivals. For completeness, we outline each of these algorithms together with the explanation of why they are not individually fair in Section 3.1. Based on the different structure of these algorithms, we employ different methods to incorporate the grace period design into them. For example, in DLP-PA, each type $i$ customer is accepted with certain probability $p_{i}$. This can lead to unfair outcomes: consider two consecutive type $i$ customers—if the first is accepted, the second will be rejected with probability $1-p_{i}$, which is the probability that the second customer perceives unfair treatment. The reverse situation also applies. When $p_{i}$ is small, the overall probability that customers experience disparate treatment increases, amplifying perceived unfairness. To address this, we avoid random decisions for individual arrivals. Instead, we divide the timeline into segments, accepting a random number of customers in a first-come-first-serve manner per segment. Grace periods inside each segment ensure fairness when decisions change from acceptance to rejection within each segment; grace periods between each segment guarantee fairness when decisions change from rejection to acceptance between the segments. For clarity, we refer to the new algorithms as GP-Enhanced.
@@
Minimal Trade-off between Revenue Maximization and Individual Fairness in an Asymptotic Regime. Contrary to the common belief that there is a fundamental trade-off between maximizing revenue and ensuring individual fairness, we show that standard RM algorithms, when augmented with the Grace Period design, can achieve fairness with minimal revenue loss in the asymptotic regime.
In addition to the fairness parameter $\alpha$ discussed above, our fairness definition (see Definition 3.1) includes a second parameter $\delta>0$, which captures the stochastic nature of the arrival process. Specifically, fairness is required to hold with probability at least $1-\delta$, allowing the algorithm to remain robust against rare, extreme arrival patterns. In the following performance summaries, we set $\delta=1/T$, where $T$ is the length of the time horizon. This choice ensures fairness holds with high probability while enabling a fair comparison across algorithms.
@@
In the stochastic arrival setting, algorithm performance is evaluated using regret, defined as the expected gap between the revenue earned by the algorithm and the revenue of the offline optimal policy with full knowledge of demand. Table 1 summarizes the regret bounds for our fairness-aware GP-Enhanced algorithms and their original counterparts. All results in the table assume $\delta=1/T$, and for any fixed $\alpha\in(0,1)$, the asymptotic regret remains unchanged—demonstrating that fairness can be achieved without sacrificing long-run revenue. For settings where the customer arrival process is adversarial, the performance of RM algorithms is measured in terms of the competitive ratio; that is, the worst-case ratio (across all possible arrival instances) between the revenue generated by the algorithm and the hindsight optimal (i.e., the maximum revenue assuming full knowledge of the realized demand). Table 2 contains a summary of the competitive ratio for GP-Enhanced algorithms for adversarial arrivals versus the traditional non-individually fair counterparts.worst-case environments.
@@
Table 1: Regret of Original Algorithms and GP-Enhanced Algorithms under Stochastic Arrivals
Original Algorithm Regret of Original Algorithm Regret of GP-Enhanced Algorithm
@@
RDLP-PA (resolve once) $O(T^{1/3})$ $O(T^{1/3})$ (Theorem 5.1)
@@
Table 2: Competitive Ratio of Original Algorithms and GP-Enhanced Algorithms under Adversarial Arrivals. $C\in[0,1]$ is a setting-specific value that is specified in the analysis, $m$ is the number of resources
. Original Algorithm Competitive Ratio of Original Algorithm Competitive Ratio of GP-Enhanced Algorithm BL $C$ $C-O(\log m/m)$ (Theorem 6.1) NESTING $C$ $C-O(\log m/m)$ (Theorem 6.2)
@@
No Additional Computational Cost for Implementing Individual Fairness. One might conjecture that simultaneously achieving optimal revenue and being individually fair might necessitate additional computational cost for the algorithm; but this is not the case. In all GP-Enhanced algorithms, we have preserved the algorithms’ original structure to the maximum extent possible, maintaining the order of computational complexity equivalent to that of the original algorithm.
@@
Survey Highlighting the Significance of Individual Fairness in RM. We conducted an online survey between October and November 2023 on Prolific about people’s perception of unfair treatment in RM. Our findings suggest: (i) When customers witness behavior in a store that is considered “unfair” (as described in Definition 3.1) before they shop, their inclination to buy decreases. (ii) A customer who has bought from a store in the past and then experiences what they perceive as unfair treatment is likely to decrease their purchasing from that store in the future. (iii) Individuals perceive different treatment happening in the near term as more unfair compared to those in the distant past. With these observations, we can evaluate the importance of individual fairness to business and validate the mathematical definition of individual fairness (Definition 3.1).
We have included a discussion on additional related work in Appendix 12, where we discuss the time variant individual fairness metrics that have been traditionally used in the literature. Importantly, we categorize most previous definitions via an axiomatic approach. We also provide a comprehensive literature review on modern revenue management models, price-based revenue management, and the group fairness in revenue management.
2 Classical Network Revenue Management Model & Preliminaries